OpenStudy (anonymous):

If z = f(x, y), where x = r2 + s2 and y = 8rs, find ∂^2z/∂r∂s. I don't get how to do second partial derivatives anyone care to show me the pattern? thanks!

OpenStudy (anonymous):

Correction: If z = f(x, y), where x = r^2 + s^2 and y = 8rs, find ∂^2z/∂r∂s. I don't get how to do second partial derivatives anyone care to show me the pattern? thanks!